OpenStudy (anonymous):

find the derivative of x^3 at x=-2 algebraically, jot my memory, do you use f(x+h)-f(x)/h?

OpenStudy (phi):

Unless you are deriving the various formulas, there is no reason to use the fundamental approach. Use the exponent rule: \[\frac{ d x^n}{dx}= n x^{n-1} \]

OpenStudy (phi):

once you have found the derivative, replace x with -2 and evaluate it to answer the question

OpenStudy (precal):

Use the power rule. Easiest rule in calculus to remember. By the way phi already stated it above.

OpenStudy (anonymous):

ok thank you soo, x^3 using the power rule, 3x^2 is the derivative, then plugging in x=-2 gives us 3(-2)^2=12?

OpenStudy (phi):

yes

OpenStudy (anonymous):

thank you phi and pre cal, im in calc 1 and it can be confusing at times, I guess the power rule is a lifesaver, much thanks

OpenStudy (precal):

use calculus notation for example y prime or f prime of x

OpenStudy (anonymous):

ok like this right f'(x)

OpenStudy (precal):

yes f'(-2)=12